Each of the following pairs consists of an electric circuit term and the corresponding mag-
netic circuit term. Which pairs are not corresponding?
(a) V and 
(b) G and 
(c)  and m
(d) IR and H
1
Expert's answer
2021-07-20T14:55:42-0400

"Relation\\
V ,\xi"

"Now\\F=q\xi\\\xi=\frac{V}{l}\\"


Both are connected

option (a) true

"G,\xi" are not connected

This is false option (b)

Now

"\xi;m"

"I=\frac{m}{V}\\\xi=\frac{V}{l}"

Are connected trur option (c)

Now

"It,H\xi"

"E=iR \\E=\xi e"

Both are connected

Option (d) is true

Option (b ) is wrong

"\xi ,G" are not relative